Systems Staff Meeting Notes 1/3/00

Systems Staff Meeting Notes
January 3, 2000

Present: Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Y2K transition went smoothly - some Administration staff got called back in and part of the UF Personnel database did not come back up (the code for it was written this Fall!). Net Services are fine.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. DLC orders to go out this week: 8 work stations, several scanners & RAM.
  2. ElasticForm order - the software has been charged to Bill's credit card, but the product has not been received yet.
  3. 60 Cows received late Thursday afternoon - currently located in the inner vault. Martha is working on a distribution list. Many of them will be destined for MSL, but probably won't be moved until the end of next week.
  4. John Van Hook received a Mac years ago that he uses at home for his Honors class. It was put on our Library inventory - library property records. John Humphrey suggested to him that he could trade in the Mac for one of our old CPU's. Bill says this will not happen.
Liaisons
  1. Liaisons weren't thrilled with having to unplug all of the work stations for Y2K.
Networking
  1. Glitch in Journalism - they were panicked about Y2K and unplugged the network CORE feed - taking them offline. We're working with NERDC to have an independent feed for us over there.
  2. Predictions are that the U.S.A. will come out ahead on of the rest of the world on Y2K compliant software.
  3. Network backups - still failing on NT8. Backup Agent has been installed on NT5. We need to put it on NT8 as well. The job is failing, although NT8 is being backed up - may do test files & restores.
  4. Suzy needs a test server. Might consider using NT7.
  5. SmathersL7 - Linux box online for forms to mail. AIX machine never did fail on Y2K!
  6. SmathersL2-6 are aliases pointing back to the proxy server.
  7. Winston put ATC data on NT5. Need to put it on NT6 as well. Somba server is not locked down well enough, although it is working. ATC data is to be used as a fallback for NOTIS failure.
  8. Resource Services timeout experiment - still in progress. People should be returning this week for further testing.
  9. Wiring - AFA to be rewired on Saturday, Jan. 8. West - estimates to Rob - for wiring phones/nets with room 219 feed is $55,000.
  10. NT2 - went down this morning and last Thursday morning. It was restarted but gave the blue screen of death and made a bad noise, although it eventually came up. It should report any problems in the event log. May be getting whacked by screech kitty attacks. Need to check to see if its patches are updated and see if we have a spare RAID drive on the shelf.
  11. Need to figure out what space is needed for server applications. Suzy would like to set up new server configuration so that the names and buildings are visible. Bill notes this is possible through the server manager. Bill wants to do a password change on all of the servers anyway, so figure out space changes and email it to Bill by the end of the week.
Software Projects Status
  1. Wiring database - has been installed on Bobbie's machine, awaiting feedback.
  2. Trouble Ticket Viewer - have received a bit of feedback having to do with the resolution setting. We should set it at 800x600, anything larger will generate a lot of service calls. Winston is working on restricting the permissions to liaisons only.
  3. OCLC tape conversions are finished.
  4. Q&A replacement database - Hilton is currently coding the first module.
  5. Debra's Trouble Ticket database is degrading - Ying made some changes to the presentation format and it seems to have helped. SQL server has problems with large amounts of BLOB data, there are some tricks for getting around it. These problems have not been duplicated on NT machines - Debra needs to try to manipulate the database on NT to see if the problems occur.
  6. Mellon II - David Allerton is almost finished with the data input.
DP Coordinator
  1. Debra has returned all of the Compustat tapes. Any patrons needing to use them should contact Peter McKay who will advise them on how to access CRSP Link via the Business College.
  2. The original CRSP tapes in the tape library belong to us. The subscription is now received on CD ROM.
  3. ICPSR is now accessed online, via the Internet. End of the tape library?
Calendar
  1. Everett out Jan. 6-12.
  2. MLK Holiday is Jan. 17.
  3. Debra out Jan. 20-27.
  4. Everett out Jan. 28-31.


Back to the Systems Homepage

Last updated January 11, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu